You're outside on a cold day (Tout = 0 °C)listening to your iPod with an external circular speaker having a diameter 0.0730 m. Because of the cold you decide to go inside where the temperature is 31.9 °C. If your favorite song begins with a 9.29-kHz tone, find the change in the diffraction angle of this tone assuming the air behaves as an ideal gas.
Answer #1

at 0 oC , the speed of sound is 331 m/s

so,

v1 / v2 = sqrt ( T1 / T2)

v / 331 = sqrt ( 31.9 + 273 / 0 + 273)

v / 331 = 1.0568

v = 349.8 m/s

so,

wavelength at  0 oC

= 331 / 9.29 * 1000 = 0.035629 m

so,

= arcsin 1.22 / d = arcsin (1.22 * 0.035629 / 0.0730) = 36.545 degree

wavelength at 31.9 oC

= 349.8 / 9.29 * 1000 = 0.03765 m

so,

= arcsin 1.22 / d = arcsin (1.22 * 0.03765 / 0.0730) = 38.996 degree

so,

change in diffraction angle, = 2.4517 degree